Suppose a hospital would like to estimate the proportion of patients who feel that physicians who care for them always communicated effectively when discussing their medical care. A pilot sample of 40 patients found that 22 reported that their physician communicated effectively. Determine the additional number of patients that need to be sampled to construct a 99​% confidence interval with a margin of error equal to 8​% to estimate this proportion.

the additional patients that need to be sampled is ____
Answer #1

Solution :

Given that,

n = 40

X = 22

\hat p = X / n = 22 / 40 = 0.55

1 - \hat p = 1 - 0.55 = 0.45

margin of error = E = 8% = 0.08

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 0.502

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p)

= (0.502/ 0.08)2 * 0.55*0.45

= 9.75

sample size = 10